3x-y=5
2x+y=15
Solve each of the following systems of equations using the method of substitution. If the system is inconcistent or dependent, say no.
I am needing a bit of refresher on this. I have been away from it too long. Can you help please. Thanks.

Answer by checkley71(8403)

ARE YOU SURE THERE ISN'T AN = SIGN IN THE SECOND EQUATION???????
3X-Y=5 OR -Y=-3X+5 OR Y=3X-5 NOW SUBSTITUTE (3X-5) FOR Y IN THE OTHER EQUATION.
2X+(3X-5)=15
2X+3X-5=15
5X=15+5
5X=20
X=20/5
X=4 ANSWER.
3*4-Y=5
12-Y=5
-Y=5-12
-Y=-7
Y=7 ANSWER.
PROOF
2*4+7=15
8+7=15
15=15